OpenStudy (anonymous):

If f(x) = x2 – 1 and g(x) = 2x – 3, what is the domain of (F*G)(X)

OpenStudy (anonymous):

Do you need steps to get to the answer

OpenStudy (anonymous):

well i worked it out and i got \[(-1,\infty)\]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[f(2x-3)=4x ^{2} -12x + 8\] (1,0) Solutions= (0,0)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

no medal or fan?

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

click best response